News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 13 January 2006 8.30am EST

During the last week the observing schedule was replanned to include an observation of GRB060108 that was accepted as a 1 day turn-around Target Of Opportunity (TOO) on Jan 10. Loads were replanned and uplinked on Jan 11. Observations of NGC 2552 and Polaris were impacted by the TOO and will be rescheduled in a later load.

Chandra passed through the first three eclipses of the 2006 winter season on January 6, 9 and 11 with nominal power and thermal performance. The remaining eclipses of the season will occur on Jan 14 and 17.

A real-time procedure was conducted on Jan 12 to disable the sun position monitor. The monitor was enabled due to the relative timing of commanding in the TOO loads and on-board commanding following the Jan 11 eclipse. In the event the timing condition arises in future schedules, adjustments will be made to ensure that the monitor is not enabled.

A Chandra press release was issued on Jan 10 describing observations of elliptical galaxies that reveal unexpected disruptive explosive activity associated with the central black holes. For details see: http://chandra.harvard.edu/press/06_releases/press_011006.html

The schedule of targets for next week is shown below and includes an observation of 3C279 coordinated with INTEGRAL.

------------------------------------------
GRB051221 ACIS-S Jan 15 Radiation Belts Jan 16 3C 279 ACIS-S/LETG Jan 17 RXC J1100.8+1033 ACIS-S Abell754-P1 ACIS-I Abell754-P4 ACIS-I PSR J0538+2817 ACIS-I Jan 18 YY Draconis ACIS-S/HETG 4U 1954+319 HRC-I Jan 19 Radiation Belts YY Draconis ACIS-S/HETG PSR J0538+2817 ACIS-I Jan 20 Sh2-271 ACIS-I Jan 21 YY Draconis ACIS-S/HETG 30 Doradus ACIS-I Radiation Belts 30 Doradus ACIS-I Jan 22

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports